if x<1, prove that (1+x)^1+x(1-x)^1-x > 1

Answer by venugopalramana(3286)

if x<1, prove that (1+x)^1+x(1-x)^1-x > 1
HAVE YOU TYPED PROPERLY?AS YOU TYPED IT BECOMES
TST
1+X+X-X^2-X>1
LHS = 1+X-X^2 =1-[X^2-2X*0.5+0.5^2]+0.5^2
=1.25-(X-0.5)^2<=1.25 SINCE (X-0.5)^2>=0
IF X<1,X-0.5<0.5....HENCE (X-0.5)^2<0.25
LHS<=1.25-0.25=1
SO PLEASE CHECK YOUR PROBLEM